# A Theory of Quantum Jumps and Its Applications to Lyman and Balmer Series

### v3

@@ 6 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to develop a theory of quantum jumps. The main advantage of this new theory is that it explicitly displays the Bohr rules for quantum jumps and it allows for the theory to be non-unitary. The solutions to the new equation are used to determine the time-scales of quantum jumps, and to calculate the radial probability density of finding the electron after its quantum jump. The obtained solutions are applied to the Lyman and Balmer series.
The theoretical results obtained in this paper are qualitatively in agreement with the experimental results given by Minev et al. [13], who showed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that takes place in a finite time. The specific values of the duration times of the considered quantum jumps in the Lyman and Balmer series are predicted to be in the range of $10^{-16}-10^{-15}$ $s$. Since the measurements in [13] were performed on the so-called artificial atoms, which are much larger than atoms, the theoretically predicted times cannot be directly compared to the experimental results; thus, they they still remain to be verfied experimentally.
In accordance with the experiment results [6-9], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital. Thus, it is proposed that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [44], or other methods [45-48] suitable to observe single orbitals in a hydrogen atom.
Acknowledgment: The author thanks Dora Musielak for valuable comments on the earlier version of this manuscript.

### v4

@@ 6 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to develop a theory of quantum jumps. The main advantage of this new theory is that it explicitly displays Bohr’s rule for quantum jumps and it allows for the theory to be non-unitary. The solutions to the new equation are used to determine the time-scales of quantum jumps, and to calculate the radial probability density of finding the electron after its quantum jump. The obtained solutions are applied to the Lyman and Balmer series.
The theoretical results obtained in this paper are qualitatively in agreement with the experimental results given by Minev et al. [13], who showed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that takes place in a finite time. The specific values of the duration times of the considered quantum jumps in the Lyman and Balmer series are predicted to be in the range of $10^{-16}-10^{-15}$ $s$. These values cannot be directly compared to the time-scales measured in [13] because they were performed on macroscopic artificial atoms. However, the time-scales predicted in this paper lie well within the range of the empirical and theoretical evaluations reported in [43-46].
In accordance with the experimental results [6-13], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital. Thus, it is suggested that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [50], or other methods [50-53] suitable to observe single orbitals in a hydrogen atom.
Acknowledgment: The author thanks two anonymous reviewers for valuable comments and suggestions that allow me to improve the original version of this paper. I am also indebted to Reviewer 2 for correcting some historical facts about quantum jumps presented in Introduction. Special thanks to Dora Musielak for reading the earlier version of this manuscript and suggesting improvements in its presentation of the results.
Funding: No funding was received for this work.
Conflict of Interest/Competing Interest: The author declares no conflict of interest.
Data availability statement: All the data generated in this research is available directly in the paper.